For the geometric sequence, what is the least value of n such that the nth term is greater than a 1000.
16, 24,36

Answer by stanbon(75887)

For the geometric sequence, what is the least value of n such that the nth term is greater than a 1000.
16, 24,36
-------------------------
a*r^(n-1) >= 1000
16*(3/2)^(n-1) >= 1000
(3/2)^(n-1) >= 62.5
---
Take the log of both sides:
(n-1)log(1.5) >= log(62.5)
---
n-1 >= 10.2
N >= 11.2
===================
Ans: n = 11
====================
Cheers,
Stan H.